the perimeter of a rectangle is 68m and length of its diagonal is 26m.find area

Grade:9

2 Answers

Vijay Mukati
8 years ago
Dear student

Hint: P = 2(L+B), D2= L2 + W2, A = LW, To find LW use (L+W)2= L2 + W2 + 2LW.

Thanks
A M S ARUN KRISHNA
8 years ago
 
P = 2(l + b)
68 = 2(l + b)
l + b = 34
d = 26
(l + b)2 = 342
l2 + b2 + 2lb = 1156
l2 + b2 = d2 (pythagoraus theorem)
l2 + b2 = 262 = 676
 676 + 2lb = 1156
2lb = 480
lb = 240
area = 240m2